# lemmasection

### v1

@@ 4 Experiments
In this section, we present a comprehensive set of experiments conducted on the PTB-XL benchmark dataset (Strodthoff et al., 2021). We first describe the datasets and the experimental settings. We then evaluate the effectiveness of the proposed post-training strategy by comparing its performance against the suggested fine-tuning strategy (Li et al., 2025) and several state-of-the-art methods, including recent ECG foundation models and task-specific approaches that have been carefully optimized. Additionally, we conduct experiments to further investigate the properties of the proposed method, including performance comparisons under varying training epochs and sample sizes, as well as an ablation study.
@@ 4.1 Datasets
The PTB-XL database is a widely used open-source dataset for evaluating ECG model performance, notable for its relatively large sample size, high-quality annotations and clearly defined usage guidelines (Wagner et al., 2020; Strodthoff, 2024). It contains 21,837 clinical 12-lead ECG recordings, each 10 seconds in duration, collected from 18,885 patients across a broad age range (0–95 years). Each recording was annotated by up to two cardiologists, with labels drawn from a set of 71 distinct ECG statements. These statements can be organized into 12 rhythm labels and 44 diagnostic labels, which can further be hierarchically grouped into 24 subclasses. Multiple labels may be associated with a single recording, reflecting the inherently multi-label nature of clinical ECG interpretation. Based on this labeling system, we consider four multi-label classification tasks: (1) all-71, which uses the complete set of 71 labels; (2) rhythm-12, which focuses on 12 rhythm categories; (3) diagnostic-44, which focuses on 44 diagnostic categories; and (4) subclass-23, which focuses on 23 hierarchical subclasses derived from the diagnostic labels. As discussed in Section 2, PTB-XL provides predefined splits into training, validation, and testing sets (approximately 8:1:1), enabling consistent and fair comparisons across methods (Strodthoff et al., 2021; Strodthoff, 2024). We follow this standard data splitting in all experiments.
@@ 4.2 Experimental Settings
In the initialization stage, we follow the training recipe recommended in Li et al. (2025), except for a shorter training length, since the largest validation macro Area Under the Receiver Operating Characteristic Curve (AUROC) is typically achieved in the early epochs. Specifically, the model is trained for 30 epochs with a batch size of 256, using AdamW with a weight decay of 0.1 and an initial learning rate of $5\times 10^{-3}$. A ReduceLROnPlateau scheduler (maximization mode) monitors macro AUROC on the validation set and reduces the learning rate by a factor of 0.1 if no improvement is observed for 10 consecutive epochs.
In the regularization stage, we also train for 30 epochs with a batch size of 256. Optimization is performed using AdamW with a weight decay of 0.05 and a learning rate of $5\times 10^{-4}$. A CosineAnnealingLR scheduler is adopted with a minimum learning rate of $10^{-5}$. In addition, we apply a droppath rate of 0.5 and a dropout rate of 0.05.
For both stages, the model is evaluated on the validation set at the end of each epoch, and the checkpoint with the highest macro AUROC is selected for final evaluation. This strategy serves as a form of early stopping regularization.
@@ 4.3 Comparison with Other Methods
This section presents a comprehensive comparison between the proposed post-training method and a range of baseline approaches, including the suggested baseline fine-tuning strategy, referred to ECGFounder (Li et al., 2025). To distinguish the post-training variant, we denote our method as ECGFounder-PT (Post-Training). Additional baselines include the foundation model HuBERT-ECG-BASE (Coppola et al., 2024) and task-specific and advanced models, such as inception1d (Ismail Fawaz et al., 2020), resnet1d101 (Strodthoff et al., 2021), resnet1d_wang (Wang et al., 2017), ensemble-7-models (Strodthoff et al., 2021), SPACETIME (Zhang et al., 2023), MSW-Transformer (Cheng et al., 2023), MULTIRESNET (Shi et al., 2023), and Chimera (Behrouz et al., 2024). The reported performance values are derived from (Strodthoff et al., 2021; Zhang et al., 2023; Cheng et al., 2023; Shi et al., 2023; Coppola et al., 2024; Behrouz et al., 2024).
@@
inception1d 0.925 0.931 0.930 0.953
xresnet1d101 0.925 0.937 0.929 0.957
resnet1d_wang 0.919 0.936 0.928 0.946
ensemble-7-models 0.929 0.939 0.933 0.965
MULTIRESNET 0.938 0.939 0.934 0.975
MSW-Transformer - 0.924 0.927 0.925
SPACETIME 0.936 0.941 0.933 0.967
Chimera 0.941 0.947 0.935 0.975
HuBERT-ECG-BASE 0.903 0.917 0.917 0.953
ECGFounder
0.915
(0.904, 0.924)
0.916
(0.902, 0.928)
0.927
(0.913, 0.939)
0.964
(0.948, 0.977)
ECGFounder-PT (ours)
0.944
(0.936, 0.951)
0.946
(0.939, 0.951)
0.938
(0.927, 0.947)
0.979
(0.965, 0.988)
Table 1: Performance comparison in terms of macro AUROC. Best and second-best results are highlighted in bold and underlined, respectively.
Table 1 presents the macro AUROC of all methods, along with 95% confidence intervals computed using 1000 bootstrapped samples. ECGFounder-PT achieves the highest AUROC on three tasks: all-71 (0.944), subclass-23 (0.938), and rhythm-12 (0.979), and the second-highest on diagnostic-44 (0.946). Compared with the original ECGFounder and HuBERT-ECG-BASE, ECGFounder-PT improves performance by 1.2%–3.3% and 2.3%-4.5%, respectively, demonstrating the effectiveness of the post-training adaptation.
In addition to foundation models, ECGFounder-PT also outperforms several recent state-of-the-art task-specific approaches. Compared to CNN-based baselines (inception1d, resnet1d101, resnet1d_wang, and ensemble-7-models) from Strodthoff et al. (2021), ECGFounder-PT yields macro AUROC gains of 0.7%–3.5%.When compared to multi-scale architectures like MULTIRESNET and MSW-Transformer, which explicitly exploit the hierarchical properties of ECG signals, it improves performance by 0.4%–5.8%. Even compared with recent state space models (SPACETIME and Chimera), ECGFounder-PT achieves consistent gains of 0.3%–1.2%, except for diagnostic-44, where Chimera slightly outperforms it by 0.1%. Collectively, these results highlight the robustness of ECGFounder-PT and its ability to bridge the adaptation gap of foundation models, achieving balanced generalization across diverse ECG classification tasks.
Although AUROC is a standard benchmark metric, it may be less sensitive in the presence of class imbalance. To address this, we additionally report macro Area Under the Precision-Recall Curve (AUPRC) in Table 2, with 95% confidence intervals computed using 1000 bootstrapped samples. Since AUPRC results for other methods are not available, the comparison is limited to ECGFounder and ECGFounder-PT. Across all tasks, ECGFounder-PT consistently outperforms ECGFounder, with relative gains ranging from 5.3% to 20.9%, underscoring the effectiveness of the post-training strategy, particularly in imbalanced data scenarios.
Method all-71 diagnostic-44 subclass-23 rhythm-12
ECGFounder
0.356
(0.349, 0.397)
0.321
(0.310, 0.358)
0.486
(0.464, 0.533)
0.604
(0.547, 0.687)
ECGFounder-PT (ours)
0.407
(0.397, 0.447)
0.388
(0.374, 0.438)
0.512
(0.480, 0.561)
0.645
(0.593, 0.739)
Table 2: Performance comparison in terms of macro AUPRC. Bold numbers indicate the best performance for each task.
4.4 Performance with Varying Training Epochs
We report the performance of the proposed post-training method (ECGFounder-PT) and the baseline fine-tuning strategy (ECGFounder) across epochs on the PTB-XL all-71 task in Figure 3. Since the validation and testing sets show similar trends, we present the results from the testing set for clarity. This analysis underscores the robustness and generalization of ECGFounder-PT compared to ECGFounder.
As shown in Figure 3(b), ECGFounder-PT consistently outperforms ECGFounder across both macro AUROC and macro AUPRC. The advantage is evident from the first epoch: ECGFounder-PT achieves a macro AUROC of 0.922 and a macro AUPRC of 0.384, compared to the baseline’s 0.762 and 0.118, respectively. This immediate superiority demonstrates the effectiveness of the post-training strategy in rapidly adapting the pre-trained model to downstream diagnostic tasks, leveraging prior knowledge more efficiently than the baseline.
The most notable difference lies in convergence dynamics and final performance. The baseline strategy exhibits a prolonged and unstable learning curve: its macro AUROC on the testing set rises gradually, peaking at 0.919 at epoch 16, but fails to stabilize, fluctuating between 0.901 and 0.919 from epochs 16 to 30. Similar oscillations are observed on the validation set, suggesting overfitting and limited generalization. In contrast, the proposed post-training strategy converges rapidly and smoothly to a higher, stable plateau. Macro AUROC surpasses 0.940 by epoch 8 and remains around 0.944 thereafter, with negligible variance. This stable convergence indicates that ECGFounder-PT effectively regularizes the optimization process, guiding the model toward a broad and generalizable optimum.
As shown in Figure 3(d), a similar trend is observed for the more challenging macro AUPRC, which is more informative for assessing performance on imbalanced datasets. The baseline strategy exhibits slow and unsteady improvement, reaching a maximum test macro AUPRC of 0.368. In contrast, ECGFounder-PT achieves a rapid and substantial gain, stabilizing around 0.435. This 18.2% relative improvement is particularly clinically meaningful, reflecting a markedly enhanced ability to identify true positives in imbalanced datasets, a common scenario in real-world ECG diagnosis.
Figure 3: Comparison of AUROC and AUPRC between the baseline training recipe and the proposed post-training strategy across epochs on PTB-XL validation and testing sets.
4.5 Performance with Varying Training Sample Sizes
Figure 4 compares the proposed post-training method (ECGFounder-PT) with the baseline fine-tuning strategy (ECGFounder) on the PTB-XL all-71 task, using varying training sample sizes from 10% to 100% of the available data. Macro AUROC and macro AUPRC are used as evaluation metrics. As in Section 4.4, since validation and test sets exhibit similar patterns, we present results from the testing set for clarity. This analysis explores the applicability of the proposed method in real-world scenarios where the availability of training data varies across medical institutions.
Overall, ECGFounder-PT consistently outperforms ECGFounder across all sample sizes, with improvements ranging from 3.0% to 9.1% in macro AUROC and 10.7% to 34.9% in macro AUPRC (Figures 4(b) and 4(d)).
The largest performance gains occur under limited data conditions. With only 10% of the training set, ECGFounder-PT achieves a macro AUROC of 0.899 compared to 0.824 for ECGFounder (+9.1%) and a macro AUPRC of 0.294 compared to 0.218 (+34.9%). These results demonstrate the substantial advantage of the post-training strategy in data-scarce environments, a common challenge in clinical practice.
As training data increases, the relative improvements decrease but remain consistent. For instance, at 30% of the training data, ECGFounder-PT achieves an AUROC of 0.925 and an AUPRC of 0.364, outperforming ECGFounder’s 0.877 and 0.301 (+5.5% and +21.0%, respectively). Even at higher data availability (e.g., 90%), ECGFounder-PT maintains an edge, reaching 0.942 AUROC and 0.402 AUPRC compared to 0.915 and 0.363, corresponding to relative gains of 2.9% and 10.7%.
These findings highlight the robustness of ECGFounder-PT: the post-training strategy yields consistent performance gains across diverse data regimes, with the most pronounced benefits under data scarcity. This property underscores its practical value for real-world deployment across diverse healthcare environments.
Figure 4: Comparison of AUROC and AUPRC between the baseline training recipe and the proposed post-training strategy across varying training sample sizes on PTB-XL validation and testing sets.
@@ 4.6 Ablation Study
To assess the contribution of each component in the proposed post-training strategy, we conducted an ablation study on the PTB-XL all-71 task. The results are reported in Table 3. In this analysis, “random initialization” refers to training the model from scratch with default weight initialization. For all other cases, the initialization was derived from ECGFounder, which was pre-trained on more than 7 million ECG recordings. “without droppath” indicates the removal of the stochastic depth operation, while “without linear probing” excludes the preview linear probing initialization. Likewise, “without dropout” and “without cosine schedule” denote the exclusion of dropout regularization and cosine annealing learning rate decay, respectively.
As shown in Table 3, every component in the proposed strategy contributes to the overall performance improvement. Among them, the stochastic depth operation and the preview linear probing initialization are the most important. Removing either of these components results in substantial performance degradation: the AUPRC decreases by approximately 4.1%–5.1% on the validation set and 0.7%–4.4% on the testing set, while AUROC decreases by 0.3%–0.4% and 1.1%–1.6% on the validation and testing sets, respectively. In contrast, the absence of dropout or cosine learning rate schedule leads to only marginal differences, suggesting that these techniques primarily provide auxiliary benefits.
It is also noteworthy that, even under random initialization, the proposed post-training strategy achieves performance comparable to the baseline strategy that leverages pre-training on more than 7 million ECGs. This demonstrates the robustness of the proposed approach and underscores its potential value for advancing ECG foundation models.
Table 3: Ablation study on the PTB-XL dataset (all-71 task). The table reports the effect of removing individual components from the proposed post-training strategy. Best and second-best results are highlighted in bold and underlined, respectively.
Method Setting Validation set Testing set
AUROC AUPRC AUROC AUPRC
Baseline Strategy random initialization 0.904 0.312 0.902 0.317
pre-training 0.924 0.348 0.915 0.356
Proposed Strategy random initialization 0.925 0.359 0.919 0.349
without droppath 0.940 0.417 0.929 0.389
without linear probing 0.939 0.413 0.934 0.404
without dropout 0.943 0.428 0.943 0.405
without cosine schedule 0.943 0.425 0.942 0.407
all components 0.943 0.435 0.944 0.407
@@ 5 Discussion
1. Performance Gains on ECG Foundation Models
@@
In this study, we evaluated the effectiveness of the proposed post-training strategy on four PTB-XL multi-label ECG classification benchmark tasks. The proposed method demonstrates consistent improvements over recent open-source ECG foundation models, confirming its ability to enhance generalization across diverse ECG classification tasks. The macro AUROC values of 0.944 (all-71), 0.946 (diagnostic-44), 0.938 (subclass-23), and 0.979 (rhythm-12) indicate 1.2%–3.3% gains over the baseline fine-tuning strategy (Li et al., 2025) and 2.3%-4.5% over another foundation model HuBERT-ECG-BASE (Coppola et al., 2024). Improvements in the macro AUPRC, ranging from 5.3% to 20.9%, compared to the baseline fine-tuning strategy, further highlight the model’s robustness in the context of imbalanced clinical datasets.
2. Comparison with State-of-the-Art
@@
It is worth noting that the proposed method consistently outperforms task-specific and advanced architectures. Baseline methods such as inception1d, xresnet1d101, resnet1d_wang, and ensemble-7-models (Strodthoff et al., 2021) exhibit competitive or even superior performance to recent ECG foundation models, likely due to their specialized design for ECG signals. Similarly, the multi-scale architecture MULTIRESNET (Shi et al., 2023) performs well, reflecting the inherent multi-scale structure in ECG data, while advanced models like SPACETIME (Zhang et al., 2023) and Chimera (Behrouz et al., 2024) highlight the significance of carefully designed architectures for ECG time-series data. These task-specific and advanced models reveal a deeper understanding of ECG signals, as they are tailored for the specific challenges of ECG classification. In contrast, foundation models like ECGFounder and HuBERT-ECG-BASE prioritize broad adaptability, which may not rely on a strong inductive bias specific to some ECG tasks. This flexibility allows foundation models to be applied across a wide range of tasks, but in certain cases, it can result in lower performance compared to specialized models, especially in the absence of an appropriate post-training strategy. As demonstrated in both our study and the work of Coppola et al. (2024), foundation models often lag behind task-specific and advanced architectures, even when pre-trained on large ECG datasets. This performance gap underscores the critical need for a robust post-training strategy. Our results, shown in Table 1, clearly illustrate that incorporating an effective post-training approach significantly enhances the performance of ECG foundation models, enabling them to match or even surpass task-specific and advanced architectures.
3. Sample Efficiency and Limited Data Scenarios
The proposed post-training strategy demonstrates substantial benefits across a wide range of sample sizes, with particularly significant improvements in scenarios where data is limited. As shown in Figure 4, when trained with only 10% of the dataset, approximately 1,700 ECG samples, the proposed method achieves a macro AUROC of 0.899 and a macro AUPRC of 0.294, compared to 0.824 and 0.218 for the baseline strategy (Li et al., 2025), resulting in relative improvements of 9.1% and 34.9%, respectively. Even with 90% of the training data, the post-training strategy continues to deliver consistent gains, with macro AUROC and AUPRC improvements of 2.9% and 10.7%, respectively. These findings highlight the effectiveness of the post-training approach in reducing the dependency on large annotated datasets, which is particularly beneficial in clinical ECG applications where labeled data is often scarce. This ability to enhance model performance with limited data underscores the practical utility of the proposed strategy in real-world medical settings, where access to extensive labeled ECG datasets may be constrained.
4. Training Stability and Convergence
A key factor driving the observed improvements is the enhanced training stability and convergence behavior facilitated by the proposed post-training strategy. Specifically, the method converges rapidly, surpassing a macro AUROC of 0.940 by epoch 8 and stabilizing at a plateau around 0.944 for the remainder of the training. In contrast, the baseline strategy exhibits slower convergence, with AUROC oscillating between 0.901 and 0.919 from epochs 16 to 30. Similarly, macro AUPRC reaches a stable value of 0.435, representing an 18.2% relative improvement over the baseline. These findings indicate that the post-training strategy effectively regularizes the optimization process, steering the model towards a more generalizable solution while mitigating overfitting. This enhanced stability not only accelerates convergence but also contributes to the overall robustness of the model, ensuring better performance across varying training conditions.
5. Ablation Study Insights
The ablation study highlights the contribution of individual components to overall performance. Stochastic depth and preview linear probing emerged as the most influential factors, with their removal leading to a decrease in macro AUPRC by 4.1%–5.1% on the validation set and 0.7%–4.4% on the testing set. The importance of stochastic depth has also been confirmed in a recent study on ECG classification (Zhou et al., 2025), underscoring its role in enhancing generalization. Preview linear probing, on the other hand, provides better initialization for the classification head, thereby boosting performance as expected. Conceptually, this strategy resonates with initialization approaches in statistical learning. For example, motivated by Zhou et al. (2013), Zhou et al. (2024) propose a sequential downsizing initialization strategy for tensor regression, which essentially performs lightweight training to induce a bias aligned with the target data. Although theoretical understanding of deep learning still lags behind its empirical success, insights from traditional statistical models may inspire more principled approaches to enhance modern deep learning methods.
6. Evaluation Metrics

### v2

@@ 4 Experiments
This section presents a comprehensive evaluation of the proposed post-training strategy. We first describe the PTB-XL benchmark dataset (Strodthoff et al., 2021), which serves as the primary evaluation platform. We then detail the experimental settings, followed by comparative analyses against the baseline fine-tuning strategies proposed in the original ECG foundation model papers (Na et al., 2024) and several state-of-the-art methods, including both foundation models and task-specific approaches.
@@ 4.1 Datasets
The PTB-XL database is a widely used open-source dataset for evaluating ECG model performance, notable for its relatively large sample size, high-quality annotations and clearly defined usage guidelines (Wagner et al., 2020; Strodthoff, 2024). It contains 21,837 clinical 12-lead ECG recordings, each 10 seconds in duration, collected from 18,885 patients across a broad age range (0-95 years). Each recording was annotated by up to two cardiologists, with labels drawn from a set of 71 distinct ECG statements. These statements can be organized into 12 rhythm labels and 44 diagnostic labels, which can further be hierarchically grouped into 24 subclasses. Multiple labels may be associated with a single recording, reflecting the inherently multi-label nature of clinical ECG interpretation. Based on this labeling system, we consider four multi-label classification tasks: (1) all-71, which uses the complete set of 71 labels; (2) rhythm-12, which focuses on 12 rhythm categories; (3) diagnostic-44, which focuses on 44 diagnostic categories; and (4) subclass-23, which focuses on 23 hierarchical subclasses derived from the diagnostic labels. As discussed in Section 2, PTB-XL provides predefined splits into training, validation, and testing sets (approximately 8:1:1), enabling consistent and fair comparisons across methods (Strodthoff et al., 2021; Strodthoff, 2024). We follow this standard data splitting in this study.
@@ 4.2 Experimental Settings
In the initialization stage, we adopt the same linear probing configuration as in (Na et al., 2024), training for 100 epochs, since the validation AUROC convergence for Transformer-based models generally requires a longer schedule. In the regularization stage, we retain the fine-tuning hyperparameters from (Na et al., 2024), but incorporate a stochastic depth rate of 0.1 and a dropout rate of 0.01 to enhance robustness and prevent overfitting. The model is evaluated on the validation set after each epoch, and the checkpoint with the highest validation macro AUROC is used for final testing. This strategy serves as a form of early stopping regularization.
@@ 4.3 Comparison with Other Methods
This section presents a comprehensive comparison between the proposed post-training method and a range of baseline approaches, including the suggested baseline fine-tuning strategy, referred to Transformer-FM (Na et al., 2024). To distinguish the post-training variant, we denote our method as Transformer-FM-PT. Additional baselines include the foundation model HuBERT-ECG-BASE (Coppola et al., 2024) and task-specific and advanced models, such as inception1d (Ismail Fawaz et al., 2020), resnet1d101 (Strodthoff et al., 2021), resnet1d_wang (Wang et al., 2017), ensemble-7-models (Strodthoff et al., 2021), SPACETIME (Zhang et al., 2023), MSW-Transformer (Cheng et al., 2023), MULTIRESNET (Shi et al., 2023), and Chimera (Behrouz et al., 2024). The reported performance values are derived from Strodthoff et al. (2021); Zhang et al. (2023); Cheng et al. (2023); Shi et al. (2023); Coppola et al. (2024); Behrouz et al. (2024).
Table 1: Performance comparison in terms of macro AUROC. The first column indicates whether a method is an ECG foundation model, and the third column specifies the main model architecture. Best results are highlighted in bold.
Foundation Model Method Architecture all-71 diagnostic-44 subclass-23 rhythm-12
No inception1d CNN 0.925 0.931 0.930 0.953
No xresnet1d101 CNN 0.925 0.937 0.929 0.957
No resnet1d_wang CNN 0.919 0.936 0.928 0.946
No ensemble-7-models Ensemble (CNN + others) 0.929 0.939 0.933 0.965
No MULTIRESNET CNN (multi-scale) 0.938 0.939 0.934 0.975
No MSW-Transformer Transformer (multi-scale) - 0.924 0.927 0.925
No SPACETIME State space model 0.936 0.941 0.933 0.967
No Chimera State space model 0.941 0.947 0.935 0.975
Yes HuBERT-ECG-BASE Transformer 0.903 0.917 0.917 0.953
Yes Transformer-FM Transformer 0.898 (0.888, 0.907) 0.870 (0.856, 0.884) 0.910 (0.896, 0.922) 0.968 (0.942, 0.984)
Yes Transformer-FM-PT Transformer 0.945 (0.938, 0.951) 0.947 (0.938, 0.954) 0.943 (0.933, 0.952) 0.975 (0.956, 0.987)
Table 1 presents the macro AUROC of all methods, along with 95% confidence intervals computed using 1000 bootstrapped samples. Transformer-FM-PT achieves the highest AUROC on four tasks: all-71 (0.945), diagnostic-44 (0.947) and subclass-23 (0.943), and rhythm-12 (0.975). Compared with the suggested fine-tuning strategy Transformer-FM, the proposed post-training strategy improves performance by 0.7%-8.9%, respectively, demonstrating the effectiveness of the post-training adaptation.
Beyond outperforming the baseline fine-tuning methods, the foundation models enhanced with our post-training strategy also surpass several recent state-of-the-art task-specific approaches. Compared to other CNN-based baselines (inception1d, resnet1d101, resnet1d_wang, and ensemble-7-models) from Strodthoff et al. (2021), Transformer-FM-PT yields macro AUROC gains of 0.9%-3.1%. When compared to multi-scale architectures like MULTIRESNET and MSW-Transformer, which explicitly exploit the hierarchical properties of ECG signals, it improves performance by 0.7%-2.5%, except for rhythm-12, where MULTIRESNET achieves the same performance. Even compared with recent state space models (SPACETIME and Chimera), Transformer-FM-PT achieves consistent gains of 0.4%-1.1%, except for diagnostic-44 and rhythm-12, where Chimera achieves the same performance. Collectively, these results highlight the robustness of post-training strategy and its ability to bridge the adaptation gap of foundation models, achieving balanced generalization across diverse ECG classification tasks.
Although AUROC is a standard benchmark metric, it may be less sensitive in the presence of class imbalance. To address this, we additionally report macro Area Under the Precision-Recall Curve (AUPRC) in Table 2, with 95% confidence intervals computed using 1000 bootstrapped samples. Since AUPRC results for other methods are not available, the comparison is limited to Transformer-FM. Across all tasks, the proposed post-training strategy consistently surpasses the baseline, achieving relative improvements of 23.3%-77.9%, underscoring the effectiveness of the post-training strategy, particularly in imbalanced data scenarios.
Table 2: Performance comparison in terms of macro AUPRC. Best results are highlighted in bold.
@@
Transformer-FM 0.307 (0.305, 0.337) 0.231 (0.224, 0.266) 0.441 (0.419, 0.488) 0.559 (0.520, 0.656)
Transformer-FM-PT 0.414 (0.405, 0.452) 0.411 (0.393, 0.453) 0.555 (0.531, 0.592) 0.689 (0.623, 0.774)
4.4 Per-Label Performance Evaluation
To gain a more detailed understanding of the effectiveness of our proposed post-training strategy, we evaluated both AUROC and AUPRC for each of the 71 ECG diagnoses on the test set. Figures 2a and 2b present the per-label AUROC and AUPRC differences between our post-training method and the baseline (Transformer-FM-PT - Transformer-PT). In these figures, bars represent the observed differences on the test set, while vertical error bars indicate the 95% confidence intervals obtained from 1,000 bootstrap resamples. For AUROC, 63 out of 71 labels exhibited positive differences, with 35 labels showing statistically significant improvements (i.e., the lower bound of the 95% confidence interval is greater than zero). Similarly, for AUPRC, 62 labels demonstrated positive differences, with 35 reaching statistical significance. For the remaining labels, most confidence intervals spanned both positive and negative values, indicating that the observed differences were not statistically significant.
Overall, these results demonstrate that our post-training strategy consistently enhances predictive performance across the majority of ECG diagnoses, providing a more granular than the macro-level metrics reported in the previous section.
Figure 2: Per-label comparison of AUROC and AUPRC differences between the proposed post-training strategy and the baseline (Transformer-FM-PT - Transformer-PT) across 71 ECG diagnoses. Bars represent the observed differences on the test set, and vertical error bars indicate the 95% confidence intervals estimated from 1,000 bootstrap resamples.
4.5 Convergence Analysis
To evaluate the convergence behavior of the proposed post-training strategy (Transformer-FM-PT) compared with the baseline (Transformer-FM), we conducted experiments on the validation set over 50 training epochs for the all-71 task. As shown in Figures 3a and 3b, both methods exhibit similar training dynamics: performance increases rapidly during the early epochs, reaches a plateau in the middle epochs, and remains relatively stable toward the end of training. Specifically, the proposed method surpasses the peak performance of the baseline within the first 4 epochs, achieving AUROC = 0.912 and AUPRC = 0.346, whereas the baseline reaches its maximum AUROC = 0.909 and AUPRC = 0.318 only later. From epoch 21 onward, the proposed method maintains AUROC values above 0.940 and AUPRC values above 0.420, indicating stable predictive accuracy in the later stages of training. Collectively, these results suggest that the proposed post-training strategy facilitates rapid attainment of strong validation performance while preserving stable optimization behavior.
Figure 3: Validation set performance of the proposed post-training strategy compared with the baseline over 50 training epochs for the all-71 task. (a) AUROC and (b) AUPRC are shown for both methods. The early stopping epoch, determined based on the validation AUROC, is highlighted with a star marker.
4.6 Data Efficiency
To evaluate the robustness and data efficiency of the proposed post-training strategy under varying amounts of training data, we trained the model using 10%, 20%, …, up to 100% of the available training samples for the all-71 task, and evaluated performance on the test set in terms of AUROC and AUPRC. As shown in Figures 4a and 4b, both metrics generally improve as the amount of training data increases and reach their highest values when the full training set is used. Notably, the proposed method achieves superior performance even with limited training data, outperforming the baseline trained on the full dataset. For example, when trained with only 30% of the available data, the proposed method attains an AUROC of 0.924 and an AUPRC of 0.370, both exceeding the baseline performance obtained using 100% of the training data (AUROC = 0.898, AUPRC = 0.307). These results indicate that the proposed post-training strategy is more data-efficient and yields improved generalization in low-data regimes, while consistently maintaining performance advantages across a wide range of training data proportions.
Figure 4: Test set performance of the proposed post-training strategy under varying proportions of training data. (a) AUROC and (b) AUPRC are reported for 10%, 20%, …, up to 100% of the training data. The dashed horizontal line indicates the baseline performance trained using 100% of the training data.
@@ 4.7 Ablation Study
To evaluate the contribution of the key components in the proposed post-training strategy, namely the preview linear probing (LP) initialization and stochastic depth (SD), we conducted an ablation study on the PTB-XL all-71 task. The results are summarized in Figure 5. For AUROC, removing the SD and LP components led to decreases of 0.8% and 3.9% on the validation set, respectively (Figure 5a). For AUPRC, the corresponding decreases were 3.5% and 25.9% (Figure 5b). These findings demonstrate that, despite the simplicity of the approach, both LP initialization and SD are crucial for enhancing post-training performance.
Figure 5: Validation set performance of the proposed post-training strategy for the ablation study on the all-71 task. (a) AUROC and (b) AUPRC are shown for different ablation settings. Bars represent performance for each setting, with color intensity reflecting the relative magnitude of each metric. Full denotes using the complete post-training strategy; w/o SD denotes removing the stochastic depth strategy; w/o LP denotes removing the linear probing initialization strategy.
@@ 5 Discussion
Performance Gains on ECG Foundation Models
@@
In this study, we evaluated the effectiveness of the proposed post-training strategy on four PTB-XL multi-label ECG classification benchmark tasks. The method consistently improves performance over recent ECG foundation models, demonstrating its ability to enhance generalization across diverse ECG classification tasks. As an example, we applied the strategy to a Transformer-based ECG foundation model. The macro AUROC increases by 0.7%-8.9% and the macro AUPRC by 23.3%-77.9% compared to the original fine-tuning strategy. Per-label analysis revealed performance gains in the majority of ECG diagnoses, with 35 diagnoses exhibiting statistically significant improvements. These results highlight the effectiveness of the proposed post-training approach in enhancing the performance of ECG foundation models.
Comparison with State-of-the-Art
@@
It is worth noting that the proposed method achieves competitive or superior performance compared with several recent state-of-the-art models across diverse ECG classification tasks. Baseline methods such as inception1d, xresnet1d101, resnet1d_wang, and ensemble-7-models (Strodthoff et al., 2021) exhibit competitive or even superior performance to recent ECG foundation models, likely due to their specialized design for ECG signals. Similarly, the multi-scale architecture MULTIRESNET (Shi et al., 2023) performs well, reflecting the inherent multi-scale structure in ECG data, while advanced models like SPACETIME (Zhang et al., 2023) and Chimera (Behrouz et al., 2024) highlight the significance of carefully designed architectures for ECG time-series data. These task-specific and advanced models reveal a deeper understanding of ECG signals, as they are tailored for the specific challenges of ECG classification. In contrast, foundation models prioritize broad adaptability, which may not rely on a strong inductive bias specific to some ECG tasks. This flexibility allows foundation models to be applied across a wide range of tasks, but in certain cases, it can result in lower performance compared to specialized models, especially in the absence of an appropriate post-training strategy. As demonstrated in both our study and the work of Coppola et al. (2024), foundation models might lag behind task-specific and advanced architectures, even when pre-trained on large ECG datasets. This performance gap underscores the critical need for a robust post-training strategy. Our results, shown in Table 1, clearly illustrate that incorporating an effective post-training approach significantly enhances the performance of ECG foundation models, enabling them to match or even surpass task-specific and advanced architectures.
Training Dynamics and Data Efficiency
The convergence and data efficiency analyses provide complementary insights into the optimization behavior and practical applicability of the proposed post-training strategy. The example ECG foundation model used in this study contains more than 90M trainable parameters, substantially exceeding the scale of commonly used task-specific ECG classification models (Strodthoff et al., 2021). While larger models offer greater representational capacity and adaptability, they also introduce increased challenges in training stability and generalization, particularly in limited-data scenarios. The proposed post-training strategy effectively addresses these challenges, enabling stable optimization and strong performance with reduced training data. Notably, the proposed method surpasses the baseline’s peak validation performance by epoch 4, and when trained with only 30% of the available data, it outperforms the baseline trained on the full dataset. These results indicate that the proposed strategy improves training dynamics and data efficiency, which is advantageous for real-world clinical applications with constrained annotation budgets.
Ablation Study Insights
This work does not aim to introduce a complex model architecture, but rather focuses on developing effective and principled training strategies. The ablation study demonstrates that removing the two proposed components—linear probing initialization and stochastic depth—leads to notable performance degradation, with decreases of 3.9% and 0.8% in macro AUROC, and 25.9% and 3.5% in macro AUPRC, respectively. These results indicate that even simple strategies can substantially enhance the performance of deep learning models when properly applied. Indeed, many fundamental components underpinning modern deep learning success, such as residual connections (He et al., 2016), batch normalization (Ioffe and Szegedy, 2015), and layer normalization (Ba et al., 2016), are conceptually simple yet highly effective. Motivated by recent studies in ECG classification (Zhou et al., 2025) and sequential downsizing initialization for tensor regression (Zhou et al., 2024), which emphasize initialization closer to target data distributions, our results suggest that domain-specific observations and insights from traditional statistical models may inspire more principled strategies for improving contemporary deep learning methods.
A General Post-training Framework
It is worth noting that the proposed post-training strategy serves as a general framework applicable to a wide range of ECG foundation models. The proposed initialization strategy for the linear head can be readily applied to almost all ECG foundation models. Similarly, the stochastic depth strategy is broadly applicable, as residual connections are fundamental components in modern neural network architectures. In this study, we use Transformer-FM as an example, primarily because Transformer are most commonly used foundation model architectures. Nevertheless, this does not imply that the proposed post-training strategy is limited to these two models; rather, it provides a generalizable framework that can be easily extended to other ECG foundation models. We encourage the research community to further explore and validate the proposed framework across different architectures and datasets.
Evaluation Metrics